"Time headway" in traffic flow is the elapsed time between the time that one car finishes passing a fixed point and the instant that the next car begins to pass that point. Let X = the time headway for two randomly chosen consecutive cars on a freeway during a period of heavy flow. The pdf of X is the following.

f(x) =

0.16e−0.16(x − 0.5)      x ≥ 0.5
0 otherwise

A)  What is the probability that time headway is at most 6 sec? (Round your answer to three decimal places.)

B) What is the probability that time headway is more than 6 sec? At least 6 sec? (Round your answers to three decimal places.)

more than 6s:

at least 6s:

C)  What is the probability that time headway is between 5 and 6 sec? (Round your answer to three decimal places.)
Answer #1

a)

P(X<6) =(1-exp(-0.16*(6-0.5)))=0.585

b)

more than 6s:=1-0.585 =0.415

at least 6s =0.415

c) probability that time headway is between 5 and 6 sec =P(5<X<6)

=exp(-0.16*(5-0.5))-exp(-0.16*(6-0.5)) =0.072